OpenStudy (kuvyojhmoob):

The volume of a conical pile of sand is increasing at a rate of 243π ft3/min, and the height of the pile always equals the radius r of the base. Express r as a function of time t (in minutes), assuming that r = 0 when t = 0.

OpenStudy (madhu.mukherjee.946):

Volume = 1/3 π r^2 h given the height = radius Volume = 1/3 π r^3 . . . Volume = 243 π * t .. in cubic feet 243 π t = 1/3 π r^3 r^3 = 729 * t r = 9 t^(1/3) feet